The Camel Company is considering two mutually exclusive projects with the following cash

flows. At what range of discount rates would you accept project A and at what range of discount

rates would you accept project B (will accept answers that are within 1% of correct answer)?

Year

Project A

Project B

0

-$800

-$1000

1

650

650

2

400

650
Answer #1

A project is accepted when its discount rate is less than or equal to its Internal rate of return (IRR).

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

A В D 1 Project B 2 Project A Year -$1,000.00 -$800.00 0 $650.00 $650.00 4 1 $400.00 $650.00 2 6 IRR 22.17% 19.43% Range of D

Cell reference -

A D 1 Project A Project B 2 Year 3 0 -800 -1000 1 650 650 400 2 650 5 =IRR(D3 : D5) =IRR(C3:C5) IRR Range of Discount rate to

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.